Which of the following planes are parallel to the line with equation "r = (3,0,1) + \lambda(3,2,-1)"

  1. "2x - 3y = 6"
  2. "x - y+z=6"
  3. "x - 27 + z = 6"
  4. "3 y + z = 6"
1
Expert's answer
2021-07-01T05:48:53-0400

A line and plane are parallell if the dot product of direction vector of the line and normal vector of the plane is zero.


Direction vector of given line is <3,2,-1>

Option1: normal vector is <2,-3,0>

<3,2,-1>.<2,-3,0>=6-6+0=0

Option2: normal vector is <1,-1,1>

<3,2,-1>.<1,-1,1>=3-2-1=0

Option2: normal vector is <1,0,1>

<3,2,-1>.<1,0,1>=3-1=2

Option2: normal vector is <0,3,1>

<3,2,-1>.<0,3,1>=6-1=5


Hence, options 1 and 2 are correct.